Consider the dissociation reaction:N2O4(g) ? 2 NO2 (g)The DGrxn?for this reaction at 55 ?C is ?0.8385 kJ/mol. The density of an equilibrium mixture of NO2and N2O4gases is found to be 5.12 g/L at 55 ?C and a certain pressure. Assuming these gases behave ideally, calculate:(a)

The degree of dissociation of N2O4.(b)

The average molecular weight of the mixture.(c)

The total pressure of the gas mixture.
Answer #1

The free enthalpy change \Delta G_{rxn} =- 0.8385 \ kJ/mol .

Temperature = T = 55 \ ^\circ C = 55 + 273 \ K = 328 \ K

R, gas constant = 8.314 \ J mol^{-1} K^{-1}

Now, \Delta \ G_{rxn} and K_{eq} are related as follows:

\Delta G_{rxn} = -RTln K_{eq}

Hence,

K_{eq} = e^{\frac{-\Delta G_{rxn}}{RT}}\\ = e^{\frac{-(-838.5 \ Jmol^{-1})}{8.314 \ J mol^{-1} K^{-1} \times 328 \ K}} \\ =e^{0.307} = 1.36

K_{eq} is also given by the following equation for our decomposition of N_2O_4_{(g)} \rightarrow 2NO_2_{(g)}

K_{eq} = \frac{[NO_2]_{eq}^2}{[N_2O_4]_{eq}}

a)

Now we can prepare a table to follow the concentration change of reaction and product through the course of the reaction.

N_2O_4_{(g)} \rightarrow 2NO_2_{(g)}

[N_2O_4] [NO_2]
Initial Concentration, mol/L 1 0
Change -x +2x
Equilibrium Concentration, mol/L 1-x 2x

Hence,

K_{eq} = \frac{[NO_2]_{eq}^2}{[N_2O_4]_{eq}} =1.36 \\ \Rightarrow \frac{(2x)^2}{1-x} =1.36 \\ \Rightarrow 4x^2 +1.36 x -1.36 = 0 \\ \Rightarrow x = 0.437 \ or \ x = -0.777

Since, x is a concentration, it cannot be negative. Hence x = 0.437.

Hence, the degree of dissociation or the fraction of moles of N_2O_4 that underwent dissociation is x = 0.437.

b)

Average molecular weight = mole fraction of N_2O_4 * molar mass of N_2O_4 + mole fraction of NO_2 * molar mass of NO_2

Molar mass of NO_2 = 46 gm/mol

Mole fraction of NO_2 in equilibrium mixture is

\frac{2x}{1-x+2x} = \frac{2x}{1+x} = \frac{2\times 0.437}{1+0.437} = \frac{0.874}{1.437} = 0.608

Molar mass of N_2O_4 = 92 gm/mol

Mole fraction of N_2O_4 in equilibrium mixture is

\frac{1-x}{1-x+2x} = \frac{1-x}{1+x} = \frac{1-0.437}{1+0.437} = \frac{0.563}{1.437} = 0.392

Hence, average molecular weight is

0.392 \times 92 \ gm/mol + 0.608 \times 46 \ gm/mol \\=36.064 \ gm/ mol + 27.968 \ gm/mol = 64.032 \ gm/mol

Hence, the average molecular weight of the reaction mixture is 64.032 gm/mol.

c)  Since the gas obeys ideal gas laws

pV = nRT for the mixture of gas.

p =\frac{nRT}{V}

Now substituting  n =\frac{ m}{Mm}

Where m = mass of the gas mixture and Mm is molar mass of the gas mixture.

p = \frac{m}{V} \times \frac{RT}{Mm}

Replace \frac{m}{V} = \rho (density).

p = \rho\times \frac{RT}{Mm}

We know,

\rho = 5.12\ gm/L \\ Mm = 60.032\ gm/mol \\ R = 0.082 \ L \ atm \ mol^{-1} K^{-1} \\ T = 328 \ K

Hence,

total pressure of the gas is

p = 5.12 \ gm/L \times \frac{0.082 \ L \ atm \ K^{-1} \ mol^{-1} \times 328 \ K }{64.032 \ gm/mol} = 2.15 \ atm

hence, the total pressure of the gas is 2.15 atm.